Grade 11Physical Chemistry

a 100 g 10% by weight urea solution is placed together with a 200 g glucose solution in a big jar and left for a long time. the vapour pressure in the jar at equilibrium ws found to be 23.6mm hg. if the vapour pressure of pure water at experimental temperature is 24mm hg, what is the mass % of glucose in its solution?

Kabir Jain
7 Years agoGrade 11

1 Answer

Arun
6 Years ago
Mass % of a compound can be shown as:
 
Mass % = (mass of solute / mass of solution) * 100
 
10% by mass urea of 100 g indicates = (100 * 10) / 100 = 10 g
 
Mass of glucose = 200 g
 
Mass % glucose = [mass of glucose / (mass of glucose + mass urea) ]* 100
 
= [200 / (200 + 10) ] * 100 = 95.21 %
 
Hence, mass percent of glucose is 95.21 %